Could you please explain the word problem for me so that I can explain to my son, In a basketball game, Bob scored 3 times as many points as Albert. The difference in their point totals was 10. How many points did each boy score? Also, this is systems of equations in algebra and he also has to label variables, write equations, and solve the system. I half way understand these problems but not fully. PLEASE HELP!!!!

Thank You
Tonya Clark

Answer by longjonsilver(2297)

First define a couple of things:
Let a = points scored by Albert
Let b = points scored by Bob

First eqn we know is that: b = 3a

Second thing we know is that b-a = 10 (since bob scored more than albert).

So, put equn1 into the second eqn, as: 3a-a = 10

ie 2a = 10
so a = 10/2
--> a = 5
hence b = 15

jon.